### v1

# Tate modules as condensed modules

Abstract.
We prove that the category of countable Tate modules over an arbitrary discrete ring embeds fully faithfully into that of condensed modules. If the base ring is of finite type, we characterize the essential image as generated by the free module of infinite countable rank under direct sums, duals and retracts. In the $\infty$-categorical context, we establish a fully faithful embedding of the $\infty$-category of countable Tate objects in perfect complexes, with uniformly bounded tor-amplitude, into the derived $\infty$-category of condensed modules. The boundedness assumption is necessary to ensure fullness, as we prove via an explicit counterexample in the unbounded case.
Introduction
Tate vector spaces were first introduced in [Lef42] (under the name of locally linearly compact topological vector spaces), and then reconsidered and developed in [BFM91] as a convenient generalization, in the context of topological vector spaces, of finite dimensional vector spaces for which some notion of dimension and determinant still make sense, and behave reasonably well. It is safe to say that the prototypical Tate $k$-vector space is the ring of Laurent series $k(\!(t)\!)$. Later [Dri06] extended this notion to families of Tate vector spaces over an affine scheme $\mathrm{Spec}\,R$, called Tate $R$-modules, and proved flat descent for the functor sending $R$ to the category of Tate $R$-modules.
The approach of [Dri06], like in the original definition, still belongs to the realm of topological linear algebra. More recently, in [BGW16], a purely categorical version (pioneered by [Bei87]) is defined and studied: the authors associate to any exact 1-category $\mathcal{C}$ another exact category $\mathsf{Tate}(\mathcal{C})$ of Tate-objects in $\mathcal{C}$, in such a way that, when $\mathcal{C}=\mathsf{Proj}^{\mathrm{fg}}_{R}$ is the category of finitely generated projective $R$-modules, $\mathsf{Tate}(\mathcal{C})$ is equivalent to the category of Tate $R$-modules defined in [Dri06]111when cardinalities restrictions are imposed, see [BGW16, Thm. 5.26] for a precise statement.. In particular, the construction of Tate objects in an exact category $\mathcal{C}$ can be iterated to obtain the exact category $n\text{-}\mathsf{Tate}(\mathcal{C})$ of (iterated) $n$-Tate objects in $\mathcal{C}$. This is particularly useful because it allows a proper setting for the so-called normally oriented (non-symmetric) tensor product between (iterated) Tate vector spaces: the normally oriented tensor product of a $n$-Tate vector space and a $m$-Tate vector space is naturally a $(n+m)$-Tate vector space (see [BGHW18]). Moreover this construction generalizes to (iterated) Tate objects in an exact category endowed with a suitable symmetric tensor product. We will come back to this point, as a motivation, later in this Introduction.
Several (derived) moduli stacks arising in formal loop spaces and representation theory for loop groups are expected to have cotangent complexes that are, when properly understood, possibly iterated Tate complexes (see, e.g. [Hen17a, Hel20]). This requires a derived and $\infty$-categorical version of iterated Tate objects that was developed in [Hen17b]. We will use this setting in the second part of our paper.
From the previous summary it is clear that the root of the theory of Tate $R$-modules has to be found in topological linear algebra. Recently, a new approach to topological algebra, vastly called Condensed Mathematics, has been proposed by D. Clausen and P. Scholze ([Sch19]). It is therefore natural to investigate the place of Tate $R$-modules (and, more generally, of Tate objects) inside Condensed Mathematics. This is what we try to do in this paper.
More precisely, we prove in Section 1 that for any commutative ring $R$ there is a fully faithful exact functor
$\mathsf{TC}_{R}\colon\mathsf{Tate}_{\aleph_{0},R}\lhook\joinrel\relbar\joinrel\rightarrow\mathsf{Cond}(\mathsf{Mod}_{R})$
where $\mathsf{Tate}_{\aleph_{0},R}:=\mathsf{Tate}_{\aleph_{0}}(\mathsf{Proj}^{\mathrm{fg}}_{R})$ is the 1-category of Tate $R$-modules whose size is countable [^2] In the topological description of [Dri06] these correspond to (linearly topologized) topological $R$-modules of the form $P\oplus Q^{\vee}$ where $P$ and $Q$ are discrete, countably generated $R$-modules, and $(-)^{\vee}$ denotes the topological linear dual. ([BGW16, Def. 5.3 and Def. 5.23]), and $\mathsf{Cond}(\mathsf{Mod}_{R})$ is the abelian category of condensed $R$-modules, i.e. the category of accessible sheaves of $R$-modules on the pro-étale site of a point ([Sch19]). We give two equivalent descriptions of the functor $\mathsf{TC}_{R}$: one topological (closer to the setting of [Dri06]) in Proposition 1.5, and the other category-theoretic (closer to [BGW16]) in Propositions 1.7 and 1.8, and we verify that $\mathsf{TC}_{R}$ actually lands in the full subcategory $\mathsf{Solid}_{R}$ of solid $R$-modules.
In Section 2 we characterize its essential image in Theorem 2.2: it is the full subcategory of $\mathsf{Solid}_{R}$ generated by the free module $\bigoplus_{\mathbb{N}}R$ of infinite countable rank under sums, duals and retracts. Although fully faithfulness of the functor $\mathsf{TC}_{R}$ holds for any ring $R$, we prove this characterization of its essential image under the assumption that $R$ is of finite type over $\mathbb{Z}$.
Section 3 is devoted to extend the fully faithfulness result to the $\infty$-category of Tate $R$-modules (as defined in [Hen17b]). This extension does not seem to be a trivial consequence of the 1-categorical result; in fact our approach to the $\infty$-categorical statement will be completely independent of the first two sections, avoiding the need for topological modules. Actually this $\infty$-categorical result turns out to hold only if we limit ourselves to perfect complexes (replacing finitely generated projective modules in the $\infty$-categorical setting) which are of fixed tor-amplitude. The result is achieved in Corollary 3.3 as a consequence of a more general embedding (see Theorem 3.2)
$\mathrm{Pro}_{\aleph_{0}}^{[a,b]}\mathopen{}\mathclose{{}\left(\mathscr{D}(R)}\right)\lhook\joinrel\relbar\joinrel\rightarrow\mathscr{D}(\mathsf{Cond}_{R})$
of Pro-complexes (which contain Tate objects in perfect complexes) with tor-amplitude in $[a,b]$ into the derived $\infty$-category $\mathscr{D}(\mathsf{Cond}_{R})$ of condensed $R$-modules. Finally, an interesting counterexample to full faithfulness for arbitrary unbounded perfect complexes is given in Section 4.
Further directions. A natural continuation of this work, presently under investigation, would consist in defining
  • a suitable first iteration $\mathsf{2-Solid}_{R}$ of the category $\mathsf{Solid}_{R}$ of solid $R$-modules;
  • a fully faithful functor
    $\mathsf{2-TC}_{R}\colon\mathsf{2-Tate}_{{\aleph}_{0}}(\mathsf{Proj}^{\mathrm{fg}}_{R})\lhook\joinrel\relbar\joinrel\rightarrow\mathsf{2-Solid}_{R}$
  • a solid version
    $-\overrightarrow{\otimes}^{\blacksquare}-\colon\mathsf{Solid}_{R}\times\mathsf{Solid}_{R}\longrightarrow\mathsf{2-Solid}_{R}$
    of the normally oriented tensor product
    $-\overrightarrow{\otimes}-\colon\mathsf{Tate}_{\aleph_{0},R}\times\mathsf{Tate}_{\aleph_{0},R}\longrightarrow\mathsf{2-Tate}_{\aleph_{0}}(\mathsf{Proj}^{\mathrm{fg}}_{R})$
    defined in [BGHW18],
in such a way that the diagram of 1-categories
${\mathsf{Tate}_{\aleph_{0},R}\times\mathsf{Tate}_{\aleph_{0},R}}$${\mathsf{Solid}_{R}\times\mathsf{Solid}_{R}}$${\mathsf{2-Tate}_{\aleph_{0}}(\mathsf{Proj}^{\mathrm{fg}}_{R})}$${\mathsf{2-Solid}_{R}}$$\scriptstyle{-\overrightarrow{\otimes}-}$$\scriptstyle{\mathsf{TC}_{R}\times\mathsf{TC}_{R}}$$\scriptstyle{-\overrightarrow{\otimes}^{\blacksquare}-}$$\scriptstyle{\mathsf{2-TC}_{R}}$
commutes. The definition of $\mathsf{2-Solid}_{R}$ has been recently given by Vieri Sbandati in his Bachelor Thesis [Sba24], where he also proves that $\mathsf{2-Solid}_{R}$ is an Abelian full subcategory of
$\mathsf{2-Cond}_{R}:=\mathsf{Cond}(\mathsf{Cond}(\mathsf{Mod}_{R}))$
stable under limits and colimits, and that this inclusion has a left adjoint (the so-called $2$-solidification functor). Once the 1-categorical situation is understood, a natural further step would be to investigate its $\infty$-categorical version.
Another interesting problem, that we were unable to solve, is to use our fully faithful functor $\mathsf{TC}_{R}$ in order to deduce Drinfeld’s flat descent of Tate $R$-modules from the known flat descent of solid $R$-modules. K. Česnavičius kindly informed us of results that go in this direction.
Finally, let us mention the recent approach of ultrasolid modules proposed in [Mar24], which provides, when working over a field, a variant of solid modules using pro-modules. Although we do not pursue this here, it would be interesting to make a precise comparison between their theory and Tate modules.
Acknowledgments. We started this project together with Andrea Maffei, and he contributed some important ideas; in particular, Section 1 should be considered as joint work with him. A first version of Proposition 1.5 was obtained by Francesco Iacca in his Master Thesis [Iac24]. We thank Chris Brav, Benjamin Hennion, Francesco Iacca for extremely useful discussions on the topics of this paper.
1. Embedding of Tate modules in condensed modules
In this section we will produce, for any commutative (discrete) ring $R$, a fully faithful exact embedding of the exact category $\mathsf{Tate}_{{\aleph}_{0},R}$ of countable Tate objects in the exact category $\mathsf{Proj}^{\mathrm{fg}}_{R}$ of finitely generated projective $R$-modules (as defined in [BGW16, Def. 5.23]), into the abelian category $\mathsf{Cond}(\mathsf{Mod}_{R})$ of condensed $R$-modules (as defined in [Sch19, Lectures 1 and 2]).
We will write $\mathsf{Cond}(\mathcal{C})$ for the category of $\mathcal{C}$-valued sheaves on the site of profinite sets (endowed with the finitary jointly surjective topology), that are left Kan extended from their restriction to $\kappa$-small profinite sets for some uncountable strong limit cardinal $\kappa$ (see [Sch19, Def. 2.11]). When $\mathcal{C}=\mathsf{Sets}$, we write $\mathsf{Cond}:=\mathsf{Cond}(\mathsf{Sets})$. When $R$ is a commutative ring and $\mathcal{C}=\mathsf{Mod}_{R}$, we will simply write $\mathsf{Cond}_{R}:=\mathsf{Cond}(\mathsf{Mod}_{R})$.
We will write $\mathsf{TopMod}_{R}$ for the category of linearly topologized topological modules over the (discrete) ring $R$, and continuous $R$-linear morphisms.
1.1. The embedding using topology
We mostly follow the conventions and notations of [BGW16], [Sch19], and [Dri06].
Definition 1.1.
Let $R$ be a commutative ring. We let
  • $\mathsf{Tate}^{\mathrm{Dr},\mathrm{el}}_{{\aleph}_{0}}(R)$ be the full subcategory of $\mathsf{TopMod}_{R}$ consisting of topological $R$-modules of the form $P\oplus Q^{\vee}$ where $P$ and $Q$ are discrete, countably generated projective $R$-modules, and $(-)^{\vee}$ denotes the topological linear dual.
  • $\mathsf{Tate}^{\mathrm{Dr}}_{{\aleph}_{0}}(R)$ be the idempotent completion of $\mathsf{Tate}^{\mathrm{Dr},\mathrm{el}}_{{\aleph}_{0}}(R)$ i.e. the full subcategory of $\mathsf{TopMod}_{R}$ of topological modules that are direct summands of objects in $\mathsf{Tate}^{\mathrm{Dr},\mathrm{el}}_{{\aleph}_{0}}(R)$ ([Dri06, Def. 3.2.1]).
  • $\mathsf{TopMod}^{\mathrm{cgwh}}_{R}$ be the full subcategory of $\mathsf{TopMod}_{R}$ consisting of topological modules whose underlying topological space is compactly generated and weakly Hausdorff.
If $\mathsf{Proj}^{\mathrm{fg}}_{R}$ is the category of finitely generated projective $R$-modules, we denote by $\mathsf{Tate}_{\aleph_{0},R}:=\mathsf{Tate}_{\aleph_{0}}(\mathsf{Proj}^{\mathrm{fg}}_{R})$ the 1-category of Tate $R$-modules whose size is countable, in the sense of [BGW16, Def. 5.3 and Def. 5.23]. Recall from [BGW16] that there is a canonical functor
$\tau\colon\mathsf{Tate}_{{\aleph}_{0},R}\longrightarrow\mathsf{TopMod}_{R}$ (1.1)
obtained by realizing formal limits and colimits of Tate objects as actual limits and colimits in topological modules.
Proposition 1.2.
Let $R$ be a commutative ring. The fully faithful inclusion $\mathsf{Tate}^{\mathrm{Dr}}_{{\aleph}_{0}}(R)\to\mathsf{TopMod}_{R}$ factors through the fully faithful inclusion $\mathsf{TopMod}^{\mathrm{cgwh}}_{R}\to\mathsf{TopMod}_{R}$, i.e. the underlying topological space of an object in $\mathsf{Tate}^{\mathrm{Dr}}_{{\aleph}_{0}}(R)$ is compactly generated and weakly Hausdorff.
Proof.
By the proof of [BGW16, Thm. 5.26], the realization functor (1.1) induces an exact equivalence of exact categories
$\mathsf{Tate}^{\mathrm{el}}_{{\aleph}_{0}}(\mathsf{Proj}^{\mathrm{fg}}_{R})\simeq\mathsf{Tate}^{\mathrm{Dr},\mathrm{el}}_{{\aleph}_{0}}(R).$
Therefore, [BGW16, Ex. 5.22] tells us that $\mathsf{Tate}^{\mathrm{Dr},\mathrm{el}}_{{\aleph}_{0}}(R)$ is a split exact category, and any of its objects is a direct summand of $R(\!(t)\!)$. Since the underlying topological space of $R(\!(t)\!)$ is metrizable [^3] Valuation theory gives the usual choice for a metric: $d(f,g):=2^{-\mathrm{ord}(f-g)}$., and any subspace of a metrizable space is metrizable, we get that any object in $\mathsf{Tate}^{\mathrm{Dr},\mathrm{el}}_{{\aleph}_{0}}(R)$ has an underlying metrizable topological space. The same argument shows that also any object in $\mathsf{Tate}^{\mathrm{Dr}}_{{\aleph}_{0}}(R)$ (which is, by definition, a direct summand of some object in $\mathsf{Tate}^{\mathrm{Dr},\mathrm{el}}_{{\aleph}_{0}}(R)$) is metrizable. We conclude since any metrizable topological space is compactly generated and (weakly) Hausdorff. ∎
Corollary 1.3.
The realization functor $\tau\colon\mathsf{Tate}_{{\aleph}_{0},R}\to\mathsf{TopMod}_{R}$ is fully faithful and factors through the inclusion $\mathsf{TopMod}^{\mathrm{cgwh}}_{R}\hookrightarrow\mathsf{TopMod}_{R}$.
Proof.
By [BGW16, Thm. 5.26], $\tau$ induces an exact equivalence between (split) exact categories
$\mathsf{Tate}_{{\aleph}_{0},R}\simeq\mathsf{Tate}^{\mathrm{Dr}}_{{\aleph}_{0}}(R),$
and we conclude by Proposition 1.2. ∎
We consider the functor [^4] This functor can be defined more generally for any T1 space, but one cannot extend it to the whole category $\mathsf{TopMod}_{R}$, as explained in [Sch19, Warning 2.14].
()¯$\underline{(-)}\colon\mathsf{TopMod}^{\mathrm{cgwh}}_{R}\longrightarrow\mathsf{Cond}_{R}\,\,\colon\,\,M\longmapsto\underline{M}:=C^{0}(\bullet,M).$
To be precise, the underlying condensed set of $\underline{M}$ is $C^{0}(-,U_{\mathrm{top}}(M))$, where $U_{\mathrm{top}}\colon\mathsf{TopMod}_{R}\to\mathsf{Top}$ is the obvious forgetful functor [^5] This would be written $C^{0}(-,U_{\mathrm{top}}(M))=\underline{U_{\mathrm{top}}(M)}$ in the notations of [Sch19], but we have not adopted this notation that would obviously be confusing here., while, for any $S\in\mathsf{ProFin}$, the $R$-module structure on $C^{0}(S,U_{\mathrm{top}}(M))$ is given by the obvious maps
[…]
where $R$ is endowed with the discrete topology.
Proposition 1.4.
The functor ()¯${\underline{(-)}}\colon\mathsf{TopMod}^{\mathrm{cgwh}}_{R}\to\mathsf{Cond}_{R}$ is fully faithful.
Proof.
Consider the functor
$C^{0}(-,\bullet)\colon\mathsf{Top}^{\mathrm{cgwh}}\longrightarrow\mathsf{Cond}\,\,\colon\,\,T\mapsto C^{0}(-,T),$
and the commutative diagram
[…]
Recall from [Sch19, Prop. 1.7, Th. 2.16] that the functor $C^{0}(-,\bullet)$ is fully faithful and factors through the full subcategory $\mathsf{qsCond}\subset\mathsf{Cond}$ of quasi-separated condensed sets, and that its corestriction admits a left adjoint $\mathsf{qsCond}\to\mathsf{Top}^{\mathrm{cgwh}}_{R}$ given by $X\mapsto X(*)_{\mathrm{top}}$. In particular, given $M,N\in\mathsf{TopMod}^{\mathrm{cgwh}}_{R}$, the functor $C^{0}(-,\bullet)$ induces a bijection
$\mathrm{Hom}_{\mathsf{Top}^{\mathrm{cgwh}}}(U_{\mathrm{top}}(M),U_{\mathrm{top}}(N))\cong\mathrm{Hom}_{\mathsf{Cond}}(U_{\mathrm{cond}}(\underline{M}),U_{\mathrm{cond}}(\underline{N}))$ (1.3)
with inverse $(-)(*)_{\mathrm{top}}$. Since $U_{\mathrm{Top}}$, $U_{\mathrm{Cond}}$ are faithful, commutativity of diagram (1.2) immediately implies that $\underline{(-)}$ is faithful. To prove that it is also full, consider a morphism $f\colon\underline{M}\to\underline{N}$ in $\mathsf{Cond}_{R}$; we will show that it is of the form $\underline{g}$ for some map $g\colon M\to N$ in $\mathsf{TopMod}^{\mathrm{cgwh}}$. Using bijection (1.3) and commutative diagram (1.2) it suffices to prove that the continuous map $(U_{\mathrm{Cond}}f)(*)_{\mathrm{top}}\colon U_{\mathrm{top}}(M)\to U_{\mathrm{top}}(N)$ induced by $f$ is a morphism of topological $R$-modules, for then we could choose $g$ to be this morphism. Forgetting the topologies, we see that the map of sets $(U_{\mathrm{Cond}}f)(*)\colon U_{\mathrm{Set}}(M)\to U_{\mathrm{Set}}(N)$ is the one underlying the $R$-modules morphism $f(*)\colon\underline{M}(*)\to\underline{N}(*)$, which gives the desired result. ∎
The previous two results immediately imply the following proposition.
Proposition 1.5.
The composite functor
[…]
is fully faithful.
1.2. A non-topological description of the embedding
We may give a reformulation of the functor $\mathrm{TC}_{R}$ of Prop. 1.5 that avoids going through topological modules.
We start by denoting
$a(\mathrm{const}_{-})\colon\mathsf{Mod}_{R}\longrightarrow\mathsf{Cond}_{R}$ (1.4)
the functor sending a module $P$ to the sheaf (condensed $R$-module) associated to the constant presheaf $\mathrm{const}_{P}$ with value $P$ on any $S\in\mathsf{ProFin}$. The functor $a(\mathrm{const}_{-})$ is fully faithful and left adjoint to the functor given by evaluation of a condensed $R$-module at the singleton $*\in\mathsf{ProFin}$.
Let $(-)_{\delta}\colon\mathsf{Mod}_{R}\to\mathsf{TopMod}_{R}$ the functor sending $P$ to itself endowed with the discrete topology $P_{\delta}$ (note that $R$ is always tacitly endowed with the discrete topology, so $P_{\delta}$ is indeed a topological $R$-module). Since discrete topological spaces are compactly generated and (weakly) Hausdorff, we actually have a functor $(-)_{\delta}\colon\mathsf{Mod}_{R}\to\mathsf{TopMod}^{\mathrm{cgwh}}_{R}$ which is a left adjoint.
Lemma 1.6.
There is a natural isomorphism
$\alpha\colon a(\mathrm{const}_{-})\longrightarrow\underline{(-)}\circ(-)_{\delta}$
of functors $\mathsf{Mod}_{R}\longrightarrow\mathsf{Cond}_{R}$.
Proof.
Since $a(\mathrm{const}_{-})$ is the fully faithful left adjoint to evaluation at *, we define $\alpha$ as the adjunct of the functorial isomorphism $P\simeq\underline{(P_{\delta})}(*)=C^{0}(*,P_{\delta})$ in $\mathsf{Mod}_{R}$. Let us prove that, for any $P\in\mathsf{Mod}_{R}$, $\alpha(P)\colon a(\mathrm{const}_{P})\to\underline{P_{\delta}}$ is indeed an isomorphism in $\mathsf{Cond}_{R}$. Since $\alpha(P)$ is a morphism in $\mathsf{Cond}_{R}$, and the forgetful functor $U_{\mathrm{Cond}}\colon\mathsf{Cond}_{R}\to\mathsf{Cond}$ is conservative, it is enough to prove that $U_{\mathrm{Cond}}(\alpha(P))$ is an isomorphism in $\mathsf{Cond}$. Let $K$ be an arbitrary set, and consider $\alpha^{\mathrm{pre}}(K)\colon\mathrm{const}_{K}\to C^{0}(-,K_{\delta})$ the morphism in $\mathsf{Fun}(\mathsf{ProFin}^{op},\mathsf{Set})$, defined by the inclusion $\alpha^{\mathrm{pre}}(K)(S)\colon K\to C^{0}(S,K_{\delta})$ of constant functions into locally constant functions, for any $S\in\mathsf{ProFin}$. It will be enough to show that the morphism $\alpha^{\mathrm{pre}}(K)$ exhibits $C^{0}(-,K_{\delta})$ as the sheafification of the constant presheaf $\mathrm{const}_{K}$. In order to achieve this, let $X\in\mathsf{Cond}$, and $\varphi\colon\mathrm{const}_{K}\to X$ a morphism in $\mathsf{Fun}(\mathsf{ProFin}^{op},\mathsf{Set})$: we will produce a unique $\psi\colon C^{0}(-,K_{\delta})\to X$ such that $\psi\circ\alpha^{\mathrm{pre}}(K)=\varphi$. Let $S\in\mathsf{ProFin}$, and $f\in C^{0}(S,K_{\delta})$. For $k\in K$, let $S_{k}:=f^{-1}(k)$ which is open in $S$, and $\cup_{k\in K}S_{k}=S$. Since $S$ is quasicompact, there is a finite subset $K^{\prime}\subset K$ such that $\cup_{k^{\prime}\in K^{\prime}}S_{k^{\prime}}=S$ (a finite open subcover), and the union is disjoint. But $X$ is a sheaf, hence $X(S)=X(\coprod_{k^{\prime}\in K^{\prime}}S_{k^{\prime}})\simeq\prod_{k^{\prime}\in K^{\prime}}X(S_{k^{\prime}})$, and we define $\psi(S)\colon C^{0}(S,K_{\delta})\to X(S)$ by sending $f$ to the family $(\varphi(S_{k^{\prime}})(k))_{k^{\prime}\in K^{\prime}}$. It is then easy to show that such a $\psi$ is the unique map $\psi\colon C^{0}(-,K_{\delta})\to X$ such that $\psi\circ\alpha^{\mathrm{pre}}(K)=\varphi$.
Now we want to extend the functor $a(\mathrm{const}_{-})$ to a functor
$\mathrm{TC}^{\prime}_{R}\colon\mathsf{Tate}_{{\aleph}_{0},R}\longrightarrow\mathsf{Cond}_{R}$
defined on objects in $\mathsf{Tate}^{\textrm{el}}_{{\aleph}_{0},R}\subset\mathsf{Tate}_{{\aleph}_{0},R}$ by
$\mathsf{Tate}^{\textrm{el}}_{{\aleph}_{0},R}\ni\text{``$\mathrm{lim}_{i}\,\mathrm{colim}_{j}$''}\,P_{ij}\longmapsto\mathrm{lim}_{i}\,\mathrm{colim}_{j}\,a(\mathrm{const}_{P_{ij}})\,\in\mathsf{Cond}_{R}$ (1.5)
where the colimit and limit in the target is taken in $\mathsf{Cond}_{R}$.
The proof of the next Proposition shows how to construct $\mathrm{TC}^{\prime}_{R}$ rigorously.
Proposition 1.7.
The functor $a(\mathrm{const}_{-})\colon\mathsf{Proj}^{\mathrm{fg}}_{R}\longrightarrow\mathsf{Cond}_{R}$ extends to an exact functor
$\mathrm{TC}^{\prime}_{R}\colon\mathsf{Tate}_{{\aleph}_{0},R}\longrightarrow\mathsf{Cond}_{R}$
defined on objects as in (1.5).
Proof.
Since $\mathsf{Cond}_{R}$ is (Abelian hence) idempotent complete, it will be enough to construct an extension $\mathsf{Tate}^{\textrm{el}}_{{\aleph}_{0},R}\to\mathsf{Cond}_{R}$. By [BGW16, Thm. 5.4 and Prop. 5.20] $\mathsf{Tate}^{\textrm{el}}_{{\aleph}_{0},R}$ is the smallest subcategory of $\mathsf{Ind}^{\mathrm{a}}_{{\aleph}_{0}}\mathsf{Pro}^{\mathrm{a}}_{{\aleph}_{0}}(\mathsf{Proj}^{\mathrm{fg}}_{R})$ containing both $\mathsf{Pro}^{\mathrm{a}}_{{\aleph}_{0}}(\mathsf{Proj}^{\mathrm{fg}}_{R})$ and $\mathsf{Ind}^{\mathrm{a}}_{{\aleph}_{0}}(\mathsf{Proj}^{\mathrm{fg}}_{R})$, and closed under extensions, and it is split exact. Therefore it suffices to produce exact functors
$F^{\mathrm{Pro}}\colon\mathsf{Pro}^{\mathrm{a}}_{{\aleph}_{0}}(\mathsf{Proj}^{\mathrm{fg}}_{R})\to\mathsf{Cond}_{R}\qquad F^{\mathrm{Ind}}\colon\mathsf{Ind}^{\mathrm{a}}_{{\aleph}_{0}}(\mathsf{Proj}^{\mathrm{fg}}_{R})\to\mathsf{Cond}_{R}$
both restricting to
$a(\mathrm{const}_{-})\colon\mathsf{Proj}^{\mathrm{fg}}_{R}\longrightarrow\mathsf{Cond}_{R}$
Now, $\mathsf{Cond}_{R}$ is complete and co-complete (i.e. satisfies (AB3) and (AB3*)[Sch19, Thm. 2.2]), in particular it has $\aleph_{0}$-filtered colimits and $\aleph_{0}$-cofiltered limits, so by the universal property of $\mathsf{Pro}^{\mathrm{a}}_{{\aleph}_{0}}$ and $\mathsf{Ind}^{\mathrm{a}}_{{\aleph}_{0}}$ we get $F^{\mathrm{Pro}}$, and $F^{\mathrm{Ind}}$ with the desired properties. Since the exact structure on $\mathsf{Tate}_{\aleph_{0}}(\mathsf{Proj}^{\mathrm{fg}})$ is $\mathsf{Tate}_{\aleph_{0}}(\mathsf{Ex^{\mathsf{Proj}^{\mathrm{fg}}_{R}}})$, and the exact structure $\mathsf{Ex}^{\mathsf{Proj}^{\mathrm{fg}}_{R}}$ in $\mathsf{Proj}^{\mathrm{fg}}_{R}$ is split, $\mathrm{TC}^{\prime}_{R}$ is, by definition, exact. ∎
Proposition 1.8.
There is an isomorphism $\theta\colon\mathrm{TC}_{R}\simeq\mathrm{TC}^{\prime}_{R}$ of functors $\mathsf{Tate}_{{\aleph}_{0},R}\longrightarrow\mathsf{Cond}_{R}$. In particular, $\mathrm{TC}^{\prime}_{R}$ is fully faithful since $\mathrm{TC}_{R}$ is (Prop. 1.5), and both are exact, since $\mathrm{TC}^{\prime}_{R}$ is (Prop. 1.7).
Proof.
It is enough to define a natural isomorphism
$\theta\,|_{\mathsf{Tate}^{\mathrm{el}}_{{\aleph}_{0},R}}\colon\mathrm{TC}^{\prime}_{R}\,|_{\mathsf{Tate}^{\mathrm{el}}_{{\aleph}_{0},R}}\longrightarrow\mathrm{TC}_{R}\,|_{\mathsf{Tate}^{\mathrm{el}}_{{\aleph}_{0},R}}$
[…]
$\mathrm{TC}^{\prime}_{R}(\text{``$\mathrm{lim}_{i}\,\mathrm{colim}_{j}$''}\,P_{ij})$ $=\mathrm{lim}_{i}\,\mathrm{colim}_{j}\,a(\mathrm{const}_{P_{ij}})$
$\simeq\mathrm{lim}_{i}\,a(\mathrm{const}_{\mathrm{colim}_{j}\,P_{ij}})$
$\simeq\mathrm{lim}_{i}\,\underline{(\mathrm{colim}_{j}\,P_{ij})_{\delta}}$
$=\mathrm{lim}_{i}\,C^{0}(-,(\mathrm{colim}_{j}\,P_{ij})_{\delta}).$
Since the discrete topology functor $(-)_{\delta}\colon\mathsf{Mod}_{R}\to\mathsf{TopMod}^{\mathrm{cgwh}}_{R}$ is left adjoint (to the forgetful one), we finally get
$\mathrm{lim}_{i}\,C^{0}(-,(\mathrm{colim}_{j}\,P_{ij})_{\delta})\simeq\mathrm{lim}_{i}\,C^{0}(-,\mathrm{colim}_{j}\,(P_{ij})_{\delta})\simeq C^{0}(-,\mathrm{lim}_{i}\,\mathrm{colim}_{j}\,(P_{ij})_{\delta})=\mathrm{TC}_{R}(\text{``$\mathrm{lim}_{i}\,\mathrm{colim}_{j}$''}\,P_{ij}).$
Recall from [And21, Proposition 3.16] that for any ring $R$ can be endowed with an analytic ring structure $R_{\blacksquare}$ by considering the functor of measures $R_{\blacksquare}[-]$ that sends a profinite set $S\cong\lim_{i}S_{i}$ to the condensed $R$-module
$R_{\blacksquare}[S]:=\operatorname*{colim}_{R^{\prime}\subseteq R}\lim_{i}R^{\prime}[S_{i}],$
where the colimit runs over finitely generated subrings $R^{\prime}$ of $R$. The associated category of solid modules, defined in [Sch19, Proposition 7.5], will be denoted $\mathsf{Solid}_{R}$. [^6] In [Sch19], this category of solid $R$-modules is denoted $\mathsf{Mod}_{R_{\blacksquare}}^{\mathrm{cond}}$.
Since the condensed $R$-module $a(\mathrm{const}_{P})$ associated to any discrete $R$-module $P$ is solid, and since the full subcategory $\mathsf{Solid}_{R}\subset\mathsf{Cond}_{R}$ is closed under colimits and limits, we get that $\mathrm{TC}_{R}$ and $\mathrm{TC}^{\prime}_{R}$ factor via fully faithful embeddings (denoted with the same symbols)
$\mathrm{TS}_{R}\simeq\mathrm{TS}^{\prime}_{R}\colon\mathsf{Tate}_{{\aleph}_{0},R}\longrightarrow\mathsf{Solid}_{R}.$
Remark 1.9.
Alternatively, one could consider the full subcategory $\mathsf{Mod}_{R}(\mathsf{Solid}_{\mathbb{Z}})\subseteq\mathsf{Cond}_{R}$ consisting of those condensed $R$-modules whose underlying condensed abelian group is solid; in other words, this amounts to replacing $R_{\blacksquare}$ by the analytic ring $(R,\mathbb{Z})_{\blacksquare}$. The same argument as before shows that the embedding $\mathrm{TC}_{R}$ also factors through this full subcategory.
2. Essential image of the embedding over a ring of finite type
We start with a first description of the essential image of the embedding $\mathrm{TS}$, which holds for any base ring.
Proposition 2.1.
The essential image of $\mathrm{TS}_{R}$ consists of direct summands of $R(\!(t)\!)=R^{\mathbb{N}}\oplus R^{(\mathbb{N})}$.
Proof.
Observe that any object in $\mathsf{Tate}_{{\aleph}_{0},\,R}$ is a split idempotent of $R(\!(t)\!)$. Then, $\mathrm{TS}_{R}\colon\mathsf{Tate}_{{\aleph}_{0},\,R}\longrightarrow\mathsf{Solid}_{R}$ being fully faithful between idempotent complete categories, it induces an equivalence between the full subcategories of (split) idempotents of $R(\!(t)\!)$ and of $\mathrm{TS}_{R}(R(\!(t)\!))$. ∎
Until the end of this section, we will assume that the base ring $R$ is a finitely generated $\mathbb{Z}$-algebra. In this case, we will show the following characterization of the essential image of Tate modules inside solid ones.
Theorem 2.2.
For $R$ a finitely generated $\mathbb{Z}$-algebra, the essential image of the functor $\mathrm{TS}_{R}\colon\mathsf{Tate}_{{\aleph}_{0},R}\to\mathsf{Solid}_{R}$ is the smallest full subcategory of $\mathsf{Solid}_{R}$ that contains the free module $\bigoplus_{\mathbb{N}}R$ of infinite countable rank and is stable under direct sums, retracts and duals.
The proof will be given at the end of the section. Before that, we need to analyze linear duality in the categories $\mathsf{Tate}_{\aleph_{0},R}$ and $\mathsf{Solid}_{R}$, and to compare them.
2.1. Duality for Tate modules
By [BGHW18, Proposition 3.5], any exact equivalence $\Psi\colon\mathscr{C}^{\mathrm{op}}\simeq\mathscr{C}$ of an idempotent complete exact category $\mathscr{C}$ extends to an exact equivalence
$\mathsf{Tate}^{\mathrm{el}}(\mathscr{C})^{\mathrm{op}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathsf{Tate}^{\mathrm{el}}(\mathscr{C})$ (2.1)
that moreover exchanges Pro and Ind objects.
We briefly recall the construction of this duality functor. First consider the functor
$\Phi\colon\mathsf{Tate}^{\mathrm{el}}(\mathscr{C})\longrightarrow\mathsf{Pro}^{\mathrm{a}}\mathsf{Ind}^{\mathrm{a}}(\mathscr{C})$
sending an elementary Tate object $V$ to the Pro-object $L\mapsto V/L$ indexed by the poset of lattices [^7] A lattice of $V$ is an admissible monic $L\to V$ such that $L\in\mathsf{Pro}^{\mathrm{a}}(\mathscr{C})$ and $V/L\in\mathsf{Ind}^{\mathrm{a}}(\mathscr{C})$, cf [BGW16, Definition 6.1.]. $L$ of $V$. The composite
$\mathsf{Tate}^{\mathrm{el}}(\mathscr{C})^{\mathrm{op}}\longrightarrow\mathopen{}\mathclose{{}\left(\mathsf{Pro}^{\mathrm{a}}\mathsf{Ind}^{\mathrm{a}}(\mathscr{C})}\right)^{\mathrm{op}}\cong\mathsf{Ind}^{\mathrm{a}}\mathsf{Pro}^{\mathrm{a}}\mathopen{}\mathclose{{}\left(\mathscr{C}^{\mathrm{op}}}\right)$
of $\Phi^{\mathrm{op}}$ with the canonical isomorphism factors through elementary Tate objects and restricts to an equivalence
$\overline{\Phi}\colon\mathsf{Tate}^{\mathrm{el}}(\mathscr{C})^{\mathrm{op}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathsf{Tate}^{\mathrm{el}}(\mathscr{C}^{\mathrm{op}}).$
The duality functor (2.1) is then obtained as the composite of $\overline{\Phi}$ with the functor $\mathsf{Tate}^{\mathrm{el}}(\Psi)\colon\mathsf{Tate}^{\mathrm{el}}(\mathscr{C}^{\mathrm{op}})\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathsf{Tate}^{\mathrm{el}}(\mathscr{C})$ induced by $\Psi$.
Applying this construction to the linear duality functor $(\mathsf{Proj}^{\mathrm{fg}}_{R})^{\mathrm{op}}\to\mathsf{Proj}^{\mathrm{fg}}_{R}$, we obtain an exact equivalence
$(-)^{\vee}\colon\mathsf{Tate}_{{\aleph}_{0},R}^{\mathrm{op}}\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathsf{Tate}_{{\aleph}_{0},R}.$ (2.2)
By the explicit description given above, one easily sees that there are canonical isomorphisms
$\mathopen{}\mathclose{{}\left(R^{\mathbb{N}}}\right)^{\vee}\cong R^{(\mathbb{N})}\qquad\text{and}\qquad\mathopen{}\mathclose{{}\left(R^{(\mathbb{N})}}\right)^{\vee}\cong R^{\mathbb{N}}.$ (2.3)
2.2. Duality for solid modules
Recall that in this section we assume that the base ring $R$ is a finitely generated $\mathbb{Z}$-algebra. In particular, for any set $I$ the product $\prod_{I}R$ is a compact projective object of $\mathsf{Solid}_{R}$.
Lemma 2.3.
The solid modules $\prod_{\mathbb{N}}R$ and $\bigoplus_{\mathbb{N}}R$ are linear dual to one another.
Proof.
It is clear that $\mathopen{}\mathclose{{}\left(\bigoplus_{\mathbb{N}}R}\right)^{\vee}\cong\prod_{\mathbb{N}}R$. We now prove that $\mathopen{}\mathclose{{}\left(\prod_{\mathbb{N}}R}\right)^{\vee}\cong\bigoplus_{\mathbb{N}}R$. We will use that the solid modules $\prod_{I}R$ for varying sets $I$ form a family of compact projective generators of $\mathsf{Solid}_{R}$. For any such set $I$ we have bijections
$\operatorname{Hom}_{\mathsf{Solid_{R}}}\mathopen{}\mathclose{{}\left(\prod_{I}R,\mathopen{}\mathclose{{}\left(\prod_{\mathbb{N}}R}\right)^{\vee}}\right)$ $\cong\operatorname{Hom}_{\mathsf{Solid_{R}}}\mathopen{}\mathclose{{}\left(\prod_{I}R\otimes^{\blacksquare}_{R}\prod_{\mathbb{N}}R,R}\right)$
$\cong\operatorname{Hom}_{\mathsf{Solid_{R}}}\mathopen{}\mathclose{{}\left(\prod_{I\times\mathbb{N}}R,R}\right)$
$\cong\bigoplus_{\mathbb{N}}\operatorname{Hom}_{\mathsf{Solid_{R}}}\mathopen{}\mathclose{{}\left(\prod_{I}R,R}\right)$
$\cong\operatorname{Hom}_{\mathsf{Solid_{R}}}\mathopen{}\mathclose{{}\left(\prod_{I}R,\bigoplus_{\mathbb{N}}R}\right)$
where the second isomorphism follows from [Sch19, Proposition 6.3.], the third from fully faithfulness of $\mathrm{TS}$ (Proposition 1.5) and the last one from compactness of $\prod_{I}R$. ∎
2.3. Characterization of the essential image
Lemma 2.4.
Let $f$ be an endomorphism of $R^{\mathbb{N}}\oplus R^{(\mathbb{N})}$ in $\mathsf{Tate}^{\mathrm{el}}_{\aleph_{0},R}$. Then $\mathrm{TS}(f^{\vee})=\mathrm{TS}(f)^{\vee}$.
Proof.
We can decompose $f$ as a matrix $\begin{pmatrix}f_{11}&f_{12}\\ f_{21}&f_{22}\end{pmatrix}$ of morphisms $f_{ij}\colon A_{i}\to A_{j}$, where we write $A_{1}:=\prod_{\mathbb{N}}R$ and $A_{2}:=\bigoplus_{\mathbb{N}}R$. It then suffices to show that $\mathrm{TS}(f_{ij}^{\vee})=\mathrm{TS}(f_{ij})^{\vee}$ for every $i,j\in\{1,2\}$.
We start with the case of $f_{11}\in\mathrm{End}_{\mathsf{Tate}_{R}}(\prod_{\mathbb{N}}R)$. Observe that by definition of morphisms of Tate objects, there are isomorphisms of abelian groups
${\operatorname{\mathrm{End}}_{\mathsf{Tate}_{R}}\mathopen{}\mathclose{{}\left(\prod_{\mathbb{N}}R}\right)}\cong{\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}R}\cong\operatorname{\mathrm{End}}_{\mathsf{Tate}_{R}}\mathopen{}\mathclose{{}\left(\bigoplus_{\mathbb{N}}R}\right).$
Since $\mathrm{TS}$ is fully faithful, we have similar isomorphisms for endomorphisms in solid $R$-modules ${\operatorname{\mathrm{End}}_{\mathsf{Cond}_{R}}\mathopen{}\mathclose{{}\left(\prod_{\mathbb{N}}R}\right)}\cong{\prod_{\mathbb{N}}\bigoplus_{\mathbb{N}}R}\cong\operatorname{\mathrm{End}}_{\mathsf{Cond}_{R}}\mathopen{}\mathclose{{}\left(\bigoplus_{\mathbb{N}}R}\right).$ Using the isomorphisms (2.3) and the ones provided by Lemma 2.3, we need to prove that in the diagram
[…]
the outer square commutes. To do so, it suffices to show that every inner triangle commutes. We already know that the left and right triangle commutes, by construction. Now for the top (respectively the bottom) triangle, commutativily follows from inspection of the explicit construction of the duality functor (2.2) (resp. that of the proof of Lemma 2.3). This proves that $\mathrm{TS}(f_{11}^{\vee})=\mathrm{TS}(f_{11})^{\vee}$. The case of an endomorphism $f_{22}$ of $\bigoplus_{\mathbb{N}}R$ is completely similar.
We now turn to the off-diagonal morphisms, but spell out the argument only for $f_{12}\colon\prod_{\mathbb{N}}R\to\bigoplus_{\mathbb{N}}R$, the case of $f_{21}$ being similar. In that case the corresponding abelian group of morphisms can be computed as
${\operatorname{Hom}_{\mathsf{Cond}_{R}}\mathopen{}\mathclose{{}\left(\prod_{\mathbb{N}}R,\bigoplus_{\mathbb{N}}R}\right)}\stackrel{{\scriptstyle\mathrm{TC}}}{{\cong}}{\operatorname{Hom}_{\mathsf{Tate}_{R}}\mathopen{}\mathclose{{}\left(\prod_{\mathbb{N}}R,\bigoplus_{\mathbb{N}}R}\right)}\cong{\bigoplus_{\mathbb{N}}\bigoplus_{\mathbb{N}}R}.$ (2.4)
The dual morphism $f_{12}^{\vee}\colon\mathopen{}\mathclose{{}\left(\bigoplus_{\mathbb{N}}R}\right)^{\vee}\to\mathopen{}\mathclose{{}\left(\prod_{\mathbb{N}}R}\right)^{\vee}$ corresponds via the isomorphisms (2.3) to a map $\prod_{\mathbb{N}}R\to\bigoplus_{\mathbb{N}}R$, which can be seen to be $f_{12}$, using the explicit description of the linear duality functor in Tate modules. Similarly on the condensed side, tracing back through the isomorphisms (2.4) and those in the proof of Lemma 2.3, we see that $\mathrm{TS}(f_{12})^{\vee}$ can be identified with $\mathrm{TS}(f_{12})$, which gives the desired result. ∎
Proposition 2.5.
The functor $\mathrm{TS}\colon\mathsf{Tate}_{{\aleph}_{0},R}\longrightarrow\mathsf{Solid}_{R}$ preserves duals.
Proof.
Let $T$ be a countable Tate $R$-module. By [BGW16, Example 7.6.] $T$ is a direct summand of $R(\!(t)\!)$. Let $p$ be an idempotent of $R(\!(t)\!)$ with kernel $T$. Then we have isomorphisms
$\mathrm{TS}(T)^{\vee}\cong(\ker\mathrm{TS}(p))^{\vee}\cong\operatorname{\mathrm{coker}}\mathopen{}\mathclose{{}\left(\mathrm{TS}(p)^{\vee}}\right)\cong\operatorname{\mathrm{coker}}\mathopen{}\mathclose{{}\left(\mathrm{TS}(p^{\vee})}\right)\cong\mathrm{TS}(T^{\vee}),$
where the third one is given by Lemma 2.4. ∎
We can now provide the characterization of the essential image of the embedding $\mathsf{Tate}_{{\aleph}_{0},R}\hookrightarrow\mathsf{Solid}_{R}$.
Proof of Theorem 2.2.
Let $\mathcal{C}$ denote the smallest full subcategory of $\mathsf{Solid}_{R}$ that contains $\bigoplus_{\mathbb{N}}R$ and is stable under direct sums, retracts and duals.
We first prove that the essential image $\operatorname{\mathrm{im}}(\mathrm{TS})$ is contained in $\mathcal{C}$. By [BGW16, Example 7.6] every countable Tate object is a summand of $R^{\mathbb{N}}\oplus R^{(\mathbb{N})}$; since $\mathcal{C}$ is idempotent complete, it suffices to show that the latter object is in $\mathcal{C}$. By definition $\mathcal{C}$ contains $\bigoplus_{\mathbb{N}}R$, hence its dual $\prod_{\mathbb{N}}R$, therefore also their sum $\bigoplus_{\mathbb{N}}R\oplus\prod_{\mathbb{N}}R$.
We now show the converse inclusion. First, it is clear that $\operatorname{\mathrm{im}}(\mathrm{TS})$ contains $\bigoplus_{\mathbb{N}}R$ and is stable under direct sums. Since any retract is a kernel of a projector and $\mathrm{TS}$ is fully faithful (by Proposition 1.5) and exact (by Proposition 1.8), it follows that $\operatorname{\mathrm{im}}(\mathrm{TS})$ is also stable under retracts. Finally, stability under taking duals is given by Proposition 2.5. This shows that $\mathcal{C}\subseteq\operatorname{\mathrm{im}}(\mathrm{TS})$, concluding the proof. ∎
3. $\infty$-categorical embedding in the bounded case
Let $R$ be an arbitrary ring. In this section, we show that countable pro-objects in the derived $\infty$-category of a ring $R$, if uniformly bounded in the sense of tor-amplitude, embed fully faithfully into the derived $\infty$-category of condensed $R$-modules.
Definition 3.1.
We will consider the following $\infty$-categories.
  • Following [Hen17b, Definition 2.1], we let $\mathscr{T}\mathrm{ate}_{R}:=\mathscr{T}\mathrm{ate}(\mathrm{Perf}_{R})$ be the stable $\infty$-category of Tate objects in perfect complexes of $R$-modules. In this section, we will restrict our attention to its full subcategory $\mathscr{T}\mathrm{ate}_{\aleph_{0},R}$ of countable Tate objects in $\mathrm{Perf}_{R}$, by which we mean those Tate objects that can be expressed using only countable filtered and cofiltered diagrams.
  • Given integers $a$ and $b$, let $\mathscr{D}(R)^{[a,b]}$ denote the full subcategory of the derived $\infty$-category $\mathscr{D}(R)$ of $R$-modules consisting of objects with tor-amplitude in $[a,b]$, where we use cohomological grading conventions.
  • Let $\mathrm{Pro}^{[a,b]}\mathopen{}\mathclose{{}\left(\mathscr{D}(R)}\right)$ denote the full subcategory of the $\infty$-category $\mathrm{Pro}\mathopen{}\mathclose{{}\left(\mathscr{D}(R)}\right)$ consisting of Pro-objects that are uniformly of tor-amplitude in $[a,b]$, i.e. that can be written as a cofiltered diagrams taking values in $\mathscr{D}(R)^{[a,b]}$. The intersection of this $\infty$-category with $\mathscr{T}\mathrm{ate}_{\aleph_{0},R}$ will be denoted $\mathscr{T}\mathrm{ate}_{\aleph_{0},R}^{[a,b]}$.
  • Given an $\infty$-category $\mathscr{C}$ with finite limits, we let $\mathsf{Cond}(\mathscr{C})$ be the $\infty$-category of condensed objects in $\mathscr{C}$, defined as accessible $\mathscr{C}$-valued sheaves on profinite sets. For $\mathscr{C}=\mathscr{D}(R)$, note that there is an equivalence $\mathsf{Cond}(\mathscr{D}(R))\simeq\mathscr{D}(\mathsf{Cond}_{R})$.
Observe that the fully faithful functor $\mathsf{Mod}_{R}\hookrightarrow\mathsf{Cond}_{R}$ defined in (1.4) induces a fully faithful functor of stable $\infty$-categories $\mathscr{D}(R)\hookrightarrow\mathscr{D}(\mathsf{Cond}_{R})$, which is left adjoint to evaluation at $*\in\mathsf{ProFin}$. Beware that, as opposed to the previous sections, we will not distinguish between an object in $\mathscr{D}(R)$ and its image through this functor.
Using the universal property of the $\infty$-category of Pro-objects and the fact that $\mathscr{D}(\mathsf{Cond}_{R})$ is complete, we obtain a functor
$\mathrm{Pro}\mathopen{}\mathclose{{}\left(\mathscr{D}(R)}\right)\longrightarrow\mathscr{D}(\mathsf{Cond}_{R})$ (3.1)
sending a Pro-object $X\colon I\to\mathscr{D}(R)$ to its limit $\lim_{i}X_{i}$ computed in condensed objects.
We can now state the main result of this section.
Theorem 3.2.
For every integers $a,b\in\mathbb{Z}$, the restriction
$\mathrm{Pro}_{\aleph_{0}}^{[a,b]}\mathopen{}\mathclose{{}\left(\mathscr{D}(R)}\right)\longrightarrow\mathscr{D}(\mathsf{Cond}_{R})$
of the canonical functor (3.1) is fully faithful.
Corollary 3.3.
The restriction $\mathscr{T}\mathrm{ate}_{\aleph_{0}}^{[a,b]}\mathopen{}\mathclose{{}\left(\mathrm{Perf}_{R}}\right)\longrightarrow\mathscr{D}(\mathsf{Cond}_{R})$ of the functor (3.1) to Tate objects is fully faithful.
Proof of theorem 3.2.
Let $V\colon I\to\mathscr{D}(R)$ and $W\colon J\to\mathscr{D}(R)$ be countable pro-objects in $\mathscr{D}(R)$ with tor-amplitude in $[a,b]$. Using the description of the mapping spaces in pro-categories, we need to show that the canonical map
$\lim_{j\in J}\operatorname*{colim}_{i\in I^{\mathrm{op}}}\operatorname{Map}_{\mathscr{D}(R)}(V_{i},W_{j})\longrightarrow\operatorname{Map}_{\mathscr{D}(\mathsf{Cond}_{R})}\big{(}\lim_{i\in I}V_{i},\lim_{j\in J}W_{j}\big{)}$ (3.2)
is an equivalence, where the limits in the right hand side are taken in $\mathscr{D}(\mathsf{Cond}_{R})$. Using the universal property of limits, the desired result will be a consequence of the following proposition. ∎
Proposition 3.4.
Let $V\colon I\to\mathscr{D}(\mathsf{Cond}_{R})$ be a countable pro-condensed object and $W\in\mathscr{D}(R)$ be a discrete module. Assume that the following boundedness hypothesis is satisfied:
there exists $a,b\in\mathbb{Z}$ such that either $V$ factors through $\mathscr{D}(\mathsf{Cond}_{R})^{[a,b]}$ or $W$ is in $\mathscr{D}(R)^{[a,b]}$. ($\star$)
Then the canonical morphism
$\gamma\colon\operatorname*{colim}_{i\in I^{\mathrm{op}}}\operatorname{Map}_{\mathscr{D}(\mathsf{Cond}_{R})}(V_{i},W)\longrightarrow\operatorname{Map}_{\mathscr{D}(\mathsf{Cond}_{R})}\big{(}\lim_{i\in I}V_{i},W\big{)}$ (3.3)
is an equivalence.
The rest of the section in devoted to the proof of proposition 3.4. We will proceed in two steps, proving that $\gamma$ is both a monomorphism of $\infty$-groupoids (i.e. an inclusion of connected components) and an effective epimorphism (i.e. surjective on $\pi_{0}$); these two statements correspond respectively to lemmas 3.6 and 3.8.
Notation 3.5.
Given an index $i\in I$, let $p_{i}$ denote the projection $\lim\limits_{j\in I}V_{j}\to V_{i}$ and $\iota_{i}\colon\mathrm{fib}(p_{i})\to\lim\limits_{j\in I}V_{j}$ its fiber.
Lemma 3.6.
The morphism $\gamma$ from (3.3) is a monomorphism.
Proof.
We wish to prove that the non-empty fibers of $\gamma$ are contractible. Consider a morphism $f\colon V_{i_{0}}\to W$ for some $i_{0}\in I$ and let $Z$ denote the fiber of $\gamma$ at the point $\gamma(f)\simeq f\circ p_{i_{0}}$, with its natural morphism $\jmath\colon Z\to\operatorname*{colim}_{j}\operatorname{Map}(V_{j},W)$. Let $\sigma\colon K\to Z$ be a morphism from a compact $\infty$-groupoid $K\in\mathscr{S}^{\omega}$. To prove the lemma, it is sufficient to show that $\sigma$ is nullhomotopic. Since $K$ is a compact object in $\mathscr{S}$ and $I$ is filtered, the morphism $\jmath\circ\sigma\colon K\to\operatorname*{colim}_{j}\operatorname{Map}(V_{j},W)$ must factor through $\operatorname{Map}(V_{i},W)$ for some index $i\in I$. Choosing $i$ large enough, we may assume that there is a map $i_{0}\to i$ in the poset $I$, so that we have an equivalence $\gamma(f)\simeq f\circ p_{i}$. Therefore the restriction of $\gamma(f)$ along the canonical map $\iota_{i}\colon\mathrm{fib}(p_{i})\to\lim V$ is the zero morphism and thus we obtain a lift in the diagram
[…]
using that the right most square is cartesian. This implies that $\sigma$ is nullhomotopic, as desired. ∎
Remark 3.7.
Although we will not use it, we note that the proof of lemma 3.6 does not make use of the boundedness assumption ($\star$3.4).
We now establish that the comparison morphism $\gamma$ is surjective on connected components, as stated more precisely in the following result.
Lemma 3.8.
Let $f\colon\lim V\to W$ be a morphism in $\mathscr{D}(\mathsf{Cond}_{R})$, with $W$ a discrete module. Assume that the boundedness hypothesis (italic-$\star$3.4). Then $f$ factors through a projection $p_{i}\colon\lim V\to V_{i}$, for some $i\in I$.
Before proving the lemma, we introduce some convenient constructions and notations.
Notation 3.9.
For $n\in\mathbb{Z}$, we consider the functor
$\Omega^{\infty+n}:=\operatorname{Map}_{\mathscr{D}(R)}(\Sigma^{n}R,-)\colon\mathscr{D}(R)\longrightarrow\mathscr{S}$
which is right adjoint to the free functor $\Sigma^{\infty}(-)\otimes\Sigma^{n}R$. Since it commutes with limits, it induces a functor $\mathscr{D}(\mathsf{Cond}_{R})\to\mathsf{Cond}(\mathscr{S})$ sending a condensed $R$-module $A$ to the condensed pointed $\infty$-groupoid $S\mapsto\Omega^{\infty+n}(A(S))$; we still denote this new functor by $\Omega^{\infty+n}$.
Notation 3.10.
Given integers $a$ and $b$, we consider the functor
$\mathrm{U}:=\prod_{n\in[a,b]}\Omega^{\infty+n}\colon\mathscr{D}(\mathsf{Cond}_{R})\longrightarrow\mathsf{Cond}(\mathscr{S})$
leaving the dependence on $a$ and $b$ implicit in the notations.
Lemma 3.11.
If $W\in\mathsf{Cond}(\mathscr{S})$ is a discrete condensed $\infty$-groupoid, then for every profinite sets $S$ we have an equivalence
$W(S)\simeq\operatorname*{colim}_{\text{covers }\mathcal{U}\text{ of }S}W(*)^{|\mathcal{U}|}.$ (3.4)
Proof.
The constant sheaf $W$ is given by the sheafification $L\mathrm{const}_{W(*)}$ of the constant presheaf with value $W(*)\in\mathscr{C}$. Following [Lur09, Proposition 6.2.2.7.] and its proof, the sheafification functor can be constructed as a transfinite composition of the functor $(-)^{\dagger}\colon\mathscr{P}(\mathsf{ProFin})\to\mathscr{P}(\mathsf{ProFin})$ given on objects by the formula
$\mathcal{F}^{\dagger}(S)\simeq\operatorname*{colim}_{\mathcal{U}}\lim_{S^{\prime}\in\mathcal{U}}\mathcal{F}(S^{\prime})$ (3.5)
where the colimit runs over all covers $\mathcal{U}$ of $S$. More precisely, there exists a regular cardinal $\kappa$ and a transfinite sequence $\alpha\mapsto T_{\alpha}$ of endofunctors of $\mathscr{P}(\mathsf{ProFin})$, indexed by the linear poset of ordinals $\alpha\leq\kappa$, such that $L\simeq T_{\kappa}$ and all maps $T_{\alpha}\to T_{\alpha^{\prime}}$ for $\alpha<\alpha^{\prime}$ are local equivalences. The sequence can be defined using transfinite induction by $T_{0}=\mathrm{id}$, $T_{\alpha+1}=T_{\alpha}^{\dagger}$ and for any limit ordinal $\beta$, $T_{\beta}=\operatorname*{colim}_{\alpha<\beta}T_{\alpha}$. Since $(-)^{\dagger}\simeq T_{1}\to T_{\kappa}\simeq L$ is a local equivalence and the right hand side of equation (3.4) is precisely $(\mathrm{const}_{W(*)})^{\dagger}$, it suffices to show that the latter presheaf is actually a sheaf.
Let $S_{1}$ and $S_{2}$ be two extremally disconnected sets. We want to prove that the natural map
$\mathrm{const}_{W(*)}^{\dagger}(S_{1}\amalg S_{2})\longrightarrow\mathrm{const}_{W(*)}^{\dagger}(S_{1})\times\mathrm{const}_{W(*)}^{\dagger}(S_{2})$
is an equivalence. Unraveling the definitions, the above morphism corresponds to the canonical map
$\operatorname*{colim}_{\text{covers }\mathcal{U}\text{ of }S_{1}\amalg S_{2}}W(*)^{|\mathcal{U}|}\longrightarrow\mathopen{}\mathclose{{}\left(\operatorname*{colim}_{\text{covers }\mathcal{U}_{1}\text{ of }S_{1}}W(*)^{|\mathcal{U}_{1}|}}\right)\times\mathopen{}\mathclose{{}\left(\operatorname*{colim}_{\text{covers }\mathcal{U}_{2}\text{ of }S_{2}}W(*)^{|\mathcal{U}_{2}|}}\right).$
But the latter is easily seen to be an equivalence since filtered colimits commute with products and covers of $S_{1}\amalg S_{2}$ are exactly pairs of covers of $S_{1}$ and $S_{2}$. ∎
Remark 3.12 (Topological inspiration for the proof of lemma 3.8).
Before showing the lemma, let us explain a proof of the topological analog of this result, namely the fact that any map of topological $R$-modules $f\colon\lim_{i}V_{i}\to W$, from the limit of a countable sequential pro-object $V$ to a discrete module $W$, factors through one of the $V_{i}$’s.
Assume towards a contradiction that for every $i\in I$, we can pick a vector $v_{i}\in\ker(p_{i})$ that does not belong to $\ker(f)$. The sequence $(v_{i})_{i\in I}$ tends to 00 as $i\to\infty$, therefore so does its image $f(v_{i})$ in $W$. Since $W$ is discrete, this sequence must be eventually null; in particular, $f(v_{k})=0$ for some $k\in\mathbb{N}$, which then contradicts $v_{k}\notin\ker(f)$. Hence $f$ must factor through one of the $V_{i}$’s.
The actual proof in $\mathscr{D}(\mathsf{Cond}_{R})$ given below is an adaptation of this topological argument.
Proof of lemma 3.8.
Since the cofiltered category $I$ is countable, there exists a coinitial map $\omega^{\mathrm{op}}\to I$ from the opposite category of the ordinal $\omega=(0\to 1\to\dots)$. Without loss of generality, we may therefore suppose that $I\cong\omega^{\mathrm{op}}$.
Assume towards a contradiction that for every $i\in I$, the morphism $f\circ\iota_{i}$ is nonzero in $\pi_{0}\operatorname{Map}(\mathrm{fib}(p_{i}),W)$. Choose a regular cardinal $\kappa$ large enough so that all the $V_{i}$ are $\kappa$-condensed objects. The set of $\Sigma^{n}R[S]$ for varying $\kappa$-small profinite sets $S$ and $n\in\mathbb{Z}$ generate $\mathscr{D}(\mathsf{Cond}^{\kappa}_{R})$ under small colimits, therefore for every $i\in I$ we may write $\mathrm{fib}(p_{i})$ as a small colimit of such generators; since $f\circ\iota_{i}\not\simeq 0$, we can pick some $\kappa$-small non-empty profinite set $S_{i}$, some integer $n_{i}$ and a morphism $v_{i}\colon\Sigma^{n_{i}}R[S_{i}]\to\mathrm{fib}(p_{i})$ such that $f\circ\iota_{i}\circ v_{i}\not\simeq 0$.
We claim that the profinite sets $S_{i}$ can be chosen independently of $i$. To see this, consider their product $S=\prod_{i\in I}S_{i}$. Note that each projection $S\to S_{i}$ admit a section so the induced map $q_{i}\colon\Sigma^{n_{i}}R[S]\to\Sigma^{n_{i}}R[S_{i}]$ is an epimorphism. It follows that the maps $\widetilde{v}_{i}=v_{i}\circ q_{i}$ still satisfy $f\circ\iota_{i}\circ\widetilde{v}_{i}\not\simeq 0$.
By the boundedness assumption ($\star$3.4), we can pick $a,b\in\mathbb{Z}$ such that either $V$ or $W$ is of tor-dimension in $[-b,-a]$, so that $n_{i}\in[a,b]$ for every $i\in I$ (recall that we use cohomological grading conventions). We will write $v^{\prime}_{i}$ for the composite
$v^{\prime}_{i}\colon\bigoplus_{n\in[a,b]}\Sigma^{n}R[S]\xrightarrow{\mathrm{proj}_{n_{i}}}\Sigma^{n_{i}}R[S]\xrightarrow{\widetilde{v}_{i}}\mathrm{fib}(p_{i})\xrightarrow{\iota_{i}}\lim V,$
which we can view equivalently as an element $v^{\prime}_{i}\in\mathrm{U}\lim V(S):=\prod_{n\in[a,b]}\Omega^{\infty+n}\lim V(S)$. Here we made use of the functor $\mathrm{U}$ defined in (3.10), implicitly depending on the interval $[a,b]$. Note that this element $v^{\prime}_{i}$ satisfies
$f(v_{i}^{\prime})\not\simeq 0\quad\text{ in }\quad\mathrm{U}W(S).$ (3.6)
The next step is to assemble the different sections $v_{i}^{\prime}$ into a single "converging sequence", in the sense of an element $v\in\mathrm{U}\lim V(S\times\mathbb{N}^{\infty})$, where $\mathbb{N}^{\infty}$ denotes the profinite set obtained as the one-point compactification of the discrete set $\mathbb{N}$. Since $\mathrm{U}$ commutes with limits, we obtain an equivalence $\mathrm{U}\lim_{j}V_{j}\simeq\lim_{j}\mathrm{U}V_{j}$ of condensed $\infty$-groupoids. Therefore, to construct the desired sequence $v$ it suffices to produce a compatible sequence of elements $v(j)$ in $\mathrm{U}V_{j}(S\times\mathbb{N}^{\infty})$, for varying $j\in I$. We proceed as follows. By construction, we have $p_{j}v_{i}^{\prime}\simeq 0$ for $i\geq j$ so that for a fixed $j\in I$, the sequence $(p_{j}v_{i}^{\prime})_{i\in I}$ in $V_{j}(S)$ is eventually null. Using the cover
$(S\times\{0\})\amalg\dots\amalg(S\times\{j-1\})\amalg(S\times\llbracket j,\infty\rrbracket)\longrightarrow S\times\mathbb{N}^{\mathrm{\infty}}$
in $\mathsf{ProFin}$, the local sections $(p_{j}v_{0}^{\prime},\dots,p_{j}v_{j-1}^{\prime},0)$ glue into a section $v(j)\in\mathrm{U}V_{j}(S\times\mathbb{N}^{\infty})$. Observe that for any $j\geq j^{\prime}$ these sections come with equivalences $p_{jj^{\prime}}v(j^{\prime})\simeq v(j)$, where $p_{jj^{\prime}}$ denotes the natural map $V_{j}\to V_{j^{\prime}}$. More generally, the $v(j)$’s are compatible in the following sense: they are part of a natural transformation $\mathrm{const}_{S\times\mathbb{N}^{\infty}}\to\mathrm{U}V$ of functors $I\to\mathsf{Cond}(\mathscr{S})$ from the constant diagram with value the representable sheaf $S\times\mathbb{N}^{\infty}$ to the diagram $\mathrm{U}V\colon j\mapsto\mathrm{U}V_{j}$; therefore the $v(j)$’s yield an element $v\in\mathrm{U}\lim V(S\times\mathbb{N}^{\mathrm{\infty}})$.
For $x\in\mathbb{N}\cup\{\infty\}$, the inclusion of profinite sets $S\times\{x\}\to S\times\mathbb{N}^{\infty}$ induces a morphism
$\mathrm{ev}_{x}\colon\mathrm{U}\lim V(S\times\mathbb{N}^{\mathrm{\infty}})\longrightarrow\mathrm{U}\lim V(S\times\{x\})$
and let $\mathrm{ev}_{x}^{W}$ denote the analog morphism for $\mathrm{U}W$. By construction of $v$, we have equivalences $\mathrm{ev}_{i}(v)\simeq v_{i}^{\prime}$ for $i\in I$ and $\mathrm{ev}_{\infty}(v)\simeq 0$. Using the naturality of $\mathrm{U}f\colon\mathrm{U}\lim V\to\mathrm{U}W$, we obtain equivalences
$\mathrm{ev}_{\infty}(\mathrm{U}f(v))\simeq\mathrm{U}f(\mathrm{ev}^{W}_{\infty})(v)\simeq 0$ (3.7)
in $(\mathrm{U}W)(S\times\{\infty\})$. Since $W$ is a constant sheaf, so is $\mathrm{U}W$ (using that the product in $\mathrm{U}$ is finite), hence using the description of constant sheaves provided by lemma 3.11 we see that $\mathrm{U}f(v)$ can be represented as a partition $S\times\mathbb{N}^{\infty}\cong T_{1}\amalg\dots\amalg T_{n}$ labeled by elements in $\mathrm{U}W(*)$. By the equivalence (3.7), this partition can be chosen so that $T_{n}$ contains $S\times\{\infty\}$ and is labeled by $0\in W(*)$. Since $S$ is compact, the tube lemma implies that $T_{n}$ contains an open set of the form $S\times\llbracket k,\infty\rrbracket$ for some $k\in\mathbb{N}$. Therefore we obtain equivalences
$\mathrm{U}f(v_{k}^{\prime})\simeq\mathrm{U}f(\mathrm{ev}_{k}(v))\simeq\mathrm{ev}_{k}(\mathrm{U}f(v))\simeq\mathrm{ev}_{k}(0)\simeq 0,$
which contradicts (3.6); this completes the proof. ∎
Proof of Proposition 3.4.
It follows directly from lemmas 3.6 and 3.8. ∎
This concludes the proof of Theorem 3.2.
4. Counter-example to fully faithfulness in the unbounded case
In this section, we prove that the boundedness assumption in Theorem 3.2 cannot be removed.
Proposition 4.1.
The functor
$\rho\colon\mathscr{T}\mathrm{ate}_{\aleph_{0}}(\mathrm{Perf}_{R})\longrightarrow\mathscr{D}(\mathsf{Cond}_{R})$ (4.1)
is not fully faithful.
Proof.
We will prove the result by exhibiting countable Tate objects $V$ and $W$ in $\mathscr{T}\mathrm{ate}_{\aleph_{0}}(\mathrm{Perf}_{R})$ for which the canonical morphism
$\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{R}}(V,W)\longrightarrow\operatorname{Map}_{\mathscr{D}(\mathsf{Cond}_{R})}(\rho(V),\rho(W)).$ (4.2)
is not an equivalence. More precisely, we will choose $V$ and $W$ to be respectively in the subcategories $\mathrm{Pro}_{\aleph_{0}}(\mathrm{Perf}_{R})$ and $\mathrm{Ind}_{\aleph_{0}}(\mathrm{Perf}_{R})$ of $\mathscr{T}\mathrm{ate}_{\aleph_{0}}(\mathrm{Perf}_{R})$. Writing $V$ as a formal limit of perfect complexes $V_{j}$ and using the inclusions $\mathrm{Ind}_{\aleph_{0}}(\mathrm{Perf}_{R})\subseteq\mathscr{D}(R)\subseteq\mathscr{D}(\mathsf{Cond}_{R})$, we will identify the mapping spaces in (4.2) as
$\operatorname{Map}_{\mathscr{T}\mathrm{ate}_{R}}(V,W)\simeq\operatorname*{colim}_{j}\operatorname{Map}_{\mathscr{D}(R)}(V_{j},W)$
and
$\operatorname{Map}_{\mathscr{D}(\mathsf{Cond}_{R})}(\rho(V),\rho(W))\simeq\operatorname{Map}_{\mathscr{D}(\mathsf{Cond}_{R})}(\lim_{j}V_{j},W).$
To prove the result, it will then suffice to show that the comparison map
$\gamma\colon\operatorname*{colim}_{j}\operatorname{Map}_{\mathscr{D}(R)}(V_{j},W)\longrightarrow\operatorname{Map}_{\mathscr{D}(\mathsf{Cond}_{R})}(\lim_{j}V_{j},W)$ (4.3)
is not an equivalence.
We now define the counterexample. Let $I$ is the filtered poset $(0\leftarrow 1\leftarrow\dots)$ and consider the pro-object $V\colon I\to\mathscr{D}(\mathsf{Cond}_{R})$ given by $V_{j}=\prod_{0\leq n\leq j}R[n]$, where the transition maps are the obvious projections.
Claim 4.2.
The limit $\lim_{j}V_{j}\simeq\prod_{n\in\mathbb{N}}R[n]$ is a discrete object in $\mathscr{D}(\mathsf{Cond}_{R})$.
Proof of the claim.
Since the inclusion $\mathscr{D}(R)\to\mathscr{D}(\mathsf{Cond}_{R})$ of discrete objects preserves colimits, it suffices to show that the canonical map $\bigoplus_{n\in\mathbb{N}}R[n]\to\prod_{n\in\mathbb{N}}R[n]$ is an equivalence. This can be tested at the level of cohomology groups; the result then follows from the fact that $H_{n}\colon\mathscr{D}(\mathsf{Cond}_{R})\to\mathsf{Cond}_{R}$ commutes with direct sums and products for every $n$. ∎
Take $W$ to be the discrete object $\prod_{n\in\mathbb{N}}R[n]$. Observe that the identity morphism $\lim_{j}V_{j}\to W$ cannot factor through any of the $V_{j}$, as any map $V_{j}\to W$ induces the zero morphism on $H_{n}$ for $n>j$. This shows that the comparison map $\gamma$ described in (4.3) is not surjective on connected components, hence not an equivalence. ∎
Remark 4.3.
Replacing $R[n]$ by $R[-n]$ yields a similar counter-example which instead is left bounded. Therefore neither of the restrictions of the functor (4.1) to $\mathscr{T}\mathrm{ate}_{\aleph_{0}}^{[a,\infty)}(\mathrm{Perf}_{R})$ nor to $\mathscr{T}\mathrm{ate}_{\aleph_{0}}^{(-\infty,a]}(\mathrm{Perf}_{R})$ is fully faithful.